The bearing of a lighthouse from ship A is 29°NE and the bearing of that same lighthouse from ship B is 61°NW. If ship A and ship B are on the east-west line 200 km apart, how far is ship B from the lighthouse?
1
Expert's answer
2020-09-28T09:44:37-0400
"Solution"

"The\ distance\ between\ ship\ B\ and\ the\ lighthouse\ can\ be\ obtained\ as\ below.\\"

Let the distance be denoted by X which is the Adjacent side of the triangle.

Use the formula; "SOH\ CAH\ TOA, \\"

"Cos\theta=\frac{Adjecent}{Hypotenuse} \implies Cos\ 29^{\circ}\ = \frac{Adjacent}{200}\\
\implies200\ Cos\ 29^{\circ}=Adjacent\\
\therefore\ Adjacent=174.92km"

Below is an image to show the triangle.